# Aligning with Human Judgement: The Role of Pairwise Preference in Large Language Model Evaluators

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated promising capabilities as automatic evaluators in assessing the quality of generated natural language. However, LLMs still exhibit biases in evaluation and often struggle to generate coherent evaluations that align with human assessments. In this work, we first conduct a systematic study of the misalignment between LLM evaluators and human judgement, revealing that existing calibration methods aimed at mitigating biases are insufficient for effectively aligning LLM evaluators. Inspired by the use of preference data in RLHF, we formulate the evaluation as a ranking problem and introduce Pairwise-preference Search (PairS), an uncertainty-guided search method that employs LLMs to conduct pairwise comparisons and efficiently ranks candidate texts. PairS achieves state-of-the-art performance on representative evaluation tasks and demonstrates significant improvements over direct scoring. Furthermore, we provide insights into the role of pairwise preference in quantifying the transitivity of LLMs and demonstrate how PairS benefits from calibration.
@@ 1 Introduction
Large Language Models (LLMs) (Brown et al., 2020; Anil et al., 2023b; OpenAI, 2023; Anil et al., 2023a) have achieved promising instruction-following abilities, enabling an efficient adaption pipeline to solve complex natural language tasks with minimal or no labelled data (Kojima et al., 2022). This unique capability is a result of training LLMs with instruction-following data (Wei et al., 2022; Chung et al., 2022) and reinforcement learning from human feedback (RLHF) (Christiano et al., 2017; Stiennon et al., 2020). In the RLHF training paradigm, a reward model is aligned with human preferences based on ranked comparison data (Ouyang et al., 2022). It enhances LLMs’ alignment with human values, thereby generating responses that better assist humans and abide by human values.
@@
Recently, LLMs have emerged as competent alternates for evaluating various aspects of textual quality (Chen et al., 2023; Zeng et al., 2024; Zheng et al., 2024b). This forms an appealing reference-free evaluation paradigm for generative tasks while avoiding expensive human-labeling costs (Liu et al., 2023a). However, the predictions from LLM evaluators are highly sensitive to the prompt design (Zhou et al., 2023b; Zheng et al., 2024b) and even biased by multiple sources, including positional bias (Wang et al., 2023b), verbosity bias (Saito et al., 2023), and contextual bias (Zhou et al., 2023a). These biases hinder fair and trustworthy applications of LLM evaluators, resulting in inconsistency and misalignment with human judgements.
@@
To mitigate biased predictions from LLMs, calibration techniques have been developed to ensure less biased outcomes (Liu et al., 2023b; Li et al., 2023a; Zhou et al., 2023a). Motivated by this line of work, we first conduct a systematic analysis of the effectiveness of calibration techniques in aligning LLM evaluators with human judgements. We find that existing calibration methods are insufficient for aligning LLM evaluators, even provided with supervised data. We discover that misalignment in evaluation is not primarily due to biased priors over evaluation scores, but rather from misaligned evaluation standards, i.e., the likelihood of LLM evaluators. Therefore, we are motivated to explore a new evaluation paradigm with LLMs to facilitate a more aligned judgement.
@@
Figure 1: Illustration of RLHF (left), direct scoring with LLM evaluators (middle), and pairwise preference search (PairS; right). Pairwise preference data have been utilized to train the reward model to align the LLM in RLHF. Leveraging this idea, PairS reframes the traditional scoring-based evaluations into a series of pairwise comparisons between selected candidate pairs.
@@
Inspired by how LLMs are aligned through a reward model with preference data in RLHF, we argue that LLM evaluators can produce more aligned evaluations by leveraging their pairwise ranking abilities. Recent work has started exploring this intrinsic property by prompting LLMs with a pair of candidates (Wang et al., 2023b; Chen et al., 2023; Qin et al., 2023; Li et al., 2024). However, the complexity and scalability of evaluating with pairwise preference have been largely overlooked. To produce a preference ranking for evaluation, existing methods require exhaustive pairwise comparisons, as they ignore the transitivity assumption [^1] Mathematical transitivity, not linguistic transitivity, refers to the property in which if A is preferred to B and B is preferred to C, then A is preferred to C. (Qin et al., 2023). This makes the evaluation procedure intractable as the number of comparisons grows quadratically (Zheng et al., 2024b; Liusie et al., 2024).
@@
In this work, we investigate the full potential of pairwise preference in LLM evaluators and its alignment with human judgements. We propose Pairwise-preference Search (PairS), an uncertainty-guided searching method that efficiently estimates the Maximum Likelihood Estimate (MLE) of preference rankings by searching through an uncertainty-pruned pairwise comparison space. PairS is more effective and query-efficient than existing pairwise baselines while significantly outperforming traditional direct scoring evaluations. We extensively examine the applicability of PairS in representative evaluation tasks, including summarization and open-ended generation tasks. We show that PairS produces more robust and transitive evaluations compared with the greedy search baseline. Lastly, we provide further insights into how pairwise preference can be used to quantify the transitivity ability of LLM evaluators and how it benefits differently from calibration compared to direct scoring. To summarise, we provide the following contributions:
  • We present a systematic analysis of the limitations of calibration in aligning direct scoring LLM evaluators with human judgements.
@@
  • We formulate the evaluation as ranking from the perspective of transitivity and propose PairS, an uncertainty-guided pairwise preference search method that efficiently estimates MLE rankings.
  • We demonstrate that PairS attains unique scalability in aligning LLM evaluations. We also provide insights about the calibration and transitivity of pairwise preferences in LLM evaluators.
@@ 2 On the Limitations of Calibration in Aligning LLM Evaluators
Figure 2: LLM evaluations are misaligned with human judgements. The score histograms on evaluating the coherence in HANNA (Chhun et al., 2022) and SummEval (Fabbri et al., 2021). We present the scores from ground human evaluations, LLMs, and LLMs after calibrations. The histograms can be interpreted as estimated score prior distributions via marginalization.
@@
where $x$ and $y$ form an input-output pair (e.g., a source text and its summary, respectively), $I$ denotes an instruction prompt, and $s$ represents the model’s rating score in its evaluation. Similarly, we can express the human posterior predictive distribution as $p_{\mathcal{H}}(s|y;x,I)$. For simplicity, we omit the $x$ and $I$ terms hereafter. When utilizing LLMs to perform Likert scale evaluations (Likert, 1932), there is known to exist a mismatch between the score distributions of LLM ratings and human annotations (Liu et al., 2023b). We start by observing the same misalignment phenomenon, as shown in Figure 2.
Since the posterior distribution consists of two components, the likelihood, and the score preference prior, we make our hypothesis that the discrepancy between the posterior distributions of LLMs and humans mainly arises from a misalignment between their respective likelihoods. To verify this hypothesis, we conduct a preliminary experiment by calibrating the LLM’s score preference prior, $p_{\mathcal{M}}(s)$ to match the human prior, $p_{\mathcal{H}}(s)$.
Score-based LLM Evaluator Calibration. Following previous calibration methods (Zhou et al., 2023a; Jiang et al., 2023b), we first assume that the likelihood terms for both the LLM and humans are the same, $p_{\mathcal{M}}(y|s)=p_{\mathcal{H}}(y|s)$. Upon this assumption, the human posterior distribution $p_{\mathcal{H}}(s|y)$ can be expressed as:
@@
We can estimate LLM’s preference prior $p_{\mathcal{M}}(s)$ by marginalizing the model’s predictions, i.e., $p_{\mathcal{M}}(s)=\sum_{y}p_{\mathcal{M}}(s|y)p(y)$. Unlike previous work that assumes a uniform prior for $p_{\mathcal{H}}(s)$ (Zhou et al., 2023a), we directly estimate it by marginalizing human annotations to achieve a more accurate estimation. The calibration is performed by shifting the model’s predictions according to Equation 2.
@@ 3.1 Evaluation as Ranking
$\mathcal{L}_{NT}(\pi)=\prod_{i,j:\pi_{i}<\pi_{j}}P(y_{\pi_{i}}\succ y_{\pi_{j}})$ (3)
where $\mathcal{L}_{NT}$ stands for the non-transitive likelihood, [^2] This likelihood can be derived from the general learning to rank likelihood, shown in Appen. B.1. which does not make any transitive assumptions for the pairwise comparison probabilities. The MLE ranking of $\mathcal{L}_{NT}$ is also equivalent to the Kemeny-Optimal Ranking (Kemeny, 1959). However, finding the optimal ranking is challenging in two aspects:
@@
However, strict transitivity is a strong assumption. To relax the strict transitivity assumption to some extent and find a better trade-off between computational complexity and performance, we propose simplifying the non-transitive likelihood in Equation 3 by assuming compositional stochastic transitivity (Latta, 1979; Oliveira et al., 2018). The derivation of this simplified likelihood is provided in Appendix B.2:
@@ 3.2 Uncertainty-Guided Pairwise-Preference Search
Figure 4: Illustration of PairS with a beam size of 2. Heatmap values show $P(A_{i}\succ B_{j})$ by Mistral 7B, deciding the search direction. Yellow-bordered lines highlight uncertainties above $U_{h}$. Solid lines represent greedy search paths, while dashed lines denote alternative paths by other beam candidates.
The PairS algorithm is described in Algorithm 1, and a demonstration of its workings is provided in Figure 3 and Figure 4. Intuitively, while the PairS-greedy greedily chooses the preferred items between the heads of two sub-arrays, PairS-beam maintains a likelihood $\mathcal{L}_{T}$ beam that keeps track of the top-k trajectory candidates. By also considering less preferred items between the two heads, PairS explores a broader search space and potentially finds a better ranking.
@@
To represent the uncertainty between two items, we follow a commonly used approach in previous works (Wan et al., 2023) and employ entropy as the uncertainty measurement. Hence, the uncertainty between a pair of comparisons can be expressed below:
@@ 3.3 A Scaling Variant of PairS
In the first stage, we sample and rank a subset of the evaluation set, referred to as the anchor set. The size of the anchor set can be statistically determined by accommodating a tolerable error margin (Cochran, 1977) (Further details and empirical evidence in Appendix §E.2). In the second stage, we use binary search to find the positions of the remaining samples relative to the anchor set, efficiently determining their appropriate positions in the overall ranking. As depicted in Figure 5, the second stage of the scaling method results in a $O(N)$ complexity, a significant improvement over the original complexity. This two-stage scaling approach allows PairS to efficiently handle large evaluation sets while maintaining the benefits of beam search and uncertainty-based pruning. The scaling variant strikes a balance between complexity and ranking performance, making it suitable for real-world applications involving large-scale datasets.
@@ 3.4 On the Benefits of Pairwise-Preference Search
Transitivity. Both PairS-direct and PairS-beam assume transitivity to some extent, implying that an LLM with better transitivity can perform more accurate and more efficient evaluations. In other words, we expect better-performing LLMs to exhibit more transitive behaviour when making pairwise comparisons. Consequently, we propose that transitivity can serve as an effective measure of an LLM’s ability as an evaluator. By quantifying and comparing the transitivity of different LLMs, we can gain insights into their relative performance and suitability for evaluation tasks. To validate this hypothesis, we provide complementary experiments and analysis in Section 5.3.
@@
Calibration. Although pairwise preference inherently aligns LLM evaluations with human judgement, pairwise prompting can still suffer from misaligned preference priors, e.g., positional bias and contextual bias (Zhao et al., 2021; Zhou et al., 2023a). Therefore, we propose applying calibration on top of PairS to ensure debiased PairS evaluations. It is worth noting that, unlike calibrating the score-based LLM evaluators, which requires external information on the distribution of human preference prior, as shown in Section 2, pairwise comparisons can fully leverage on a uniform human prior, due to the random nature of pairwise sample selections. Hence, we implement Batch Calibration (Zhou et al., 2023a) with a uniform human prior and an estimated model prior by marginalizing LLM’s predictions over a set of comparison samples.
4 Related Works
@@
LLMs as Evaluators. With the emergent property of LLMs in following human instructions, LLMs have recently been utilized to automatically evaluate natural language generation tasks (Zhong et al., 2022; Chen et al., 2023; Wang et al., 2023a; Tan et al., 2024). LLM evaluators enable either reference-free or reference-guided metrics while being cheaper alternates to human evaluations in a diverse task set, such as summarization (Shen et al., 2023), machine translation (Kocmi & Federmann, 2023), open-ended story generation (Chiang & Lee, 2023), and instruction-following (Zeng et al., 2024). LLM evaluators are usually equipped with an evaluation criteria prompt, and the generated score in the output is then retrieved to access the quality of the candidate text (Liu et al., 2023a). This forms the scoring-based LLM evaluator that enables easy adaption to various tasks by simple instructions (Fu et al., 2023). To fairly examine various LLM evaluators, efforts have been devoted to developing meta-evaluation benchmarks (Wang et al., 2023b; Zhang et al., 2023; Zheng et al., 2024b).
@@
Pairwise Preference. Preference data have been widely used in training the reward model to reflect human preferences in RLHF (Ouyang et al., 2022). For example, Llama-2 has leveraged binary preference with ‘chosen’ or ‘rejected’ responses in its reward modelling (Touvron et al., 2023). Following the preference idea, using pairwise comparisons in evaluating the LLM outputs becomes a new paradigm (Chen et al., 2023; Shen et al., 2023), where Wang et al. (2023b) find that pairwise comparison exhibits better human correlations than traditional scoring-based LLM evaluators. However, Zheng et al. (2024b) point out the intractability of pairwise comparisons due to its quadratic query complexity (Liusie et al., 2024), which significantly limits its practicability. Qin et al. (2023) propose pairwise ranking with sorting but ignore the important transitivity of LLM predictions. In our work, we address the aforementioned issues by searching in a tractable pairwise search and transitive space based on model uncertainty.

### v3

@@ Abstract
Large Language Models (LLMs) have demonstrated promising capabilities as automatic evaluators in assessing the quality of generated natural language. However, LLMs still exhibit biases in evaluation and often struggle to generate coherent evaluations that align with human assessments. In this work, we first conduct a systematic study of the misalignment between LLM evaluators and human evaluation, revealing that existing calibration methods aimed at mitigating biases of LLMs are insufficient for effectively aligning LLM evaluators. Inspired by the use of preference data in RLHF, we formulate the evaluation as a ranking problem and introduce Pairwise-preference Search (PairS), an uncertainty-guided search method that employs LLMs to conduct pairwise comparisons locally and efficiently ranks candidate texts globally. PairS achieves state-of-the-art performance on representative evaluation tasks in long-form generations and demonstrates significant improvements over direct scoring. Furthermore, we provide insights into the role of pairwise preference in quantifying the transitivity of LLMs and demonstrate how PairS benefits from calibration using debiased pairwise evaluations.
@@ 1 Introduction
Large Language Models (LLMs) (Brown et al., 2020; OpenAI, 2023; Anil et al., 2023a; b) have achieved promising instruction-following abilities, enabling an efficient adaptation pipeline to solve complex natural language tasks with minimal or no labelled data (Kojima et al., 2022; Zhou et al., 2024c). This unique capability stems from supervised training LLMs with instruction-following data (Wei et al., 2022; Chung et al., 2022) and reinforcement learning from human feedback (RLHF) (Christiano et al., 2017; Stiennon et al., 2020). In the RLHF training paradigm, a reward model is aligned with human preferences based on ranked comparison data (Ouyang et al., 2022). It enhances LLMs’ alignment with human values (Touvron et al., 2023), thereby generating responses that better assist humans and abide by human values (Dai et al., 2024).
@@
Recently, LLMs have emerged as competent and inexpensive assessors for the evaluation of various aspects, such as coherence, fluency, consistency and hallucination, of text quality (Chen et al., 2023; Zeng et al., 2024; Zheng et al., 2024). Put simply, LLMs can be seen as an appealing reference-free evaluation paradigm for generative tasks while avoiding expensive human-labeling costs (Liu et al., 2023a). However, the predictions from LLM evaluators are highly sensitive to the prompt design (Zhou et al., 2023; Zheng et al., 2024) and even biased by multiple sources, including positional bias (Wang et al., 2023b), verbosity bias (Saito et al., 2023), and contextual bias (Zhou et al., 2024b). These biases still hinder fair and trustworthy applications of LLM evaluators, resulting in inconsistency and misalignment with human judgements.
@@
To mitigate biased predictions from LLMs, calibration techniques have been developed to ensure less biased outcomes (Liu et al., 2023b; Li et al., 2023a; Zhou et al., 2024b). Motivated by this line of work, we first conduct a systematic analysis of the effectiveness of calibration techniques in aligning LLM evaluators in direct scoring (Figure 1, middle) with human judgements. We find that existing calibration methods are insufficient for aligning LLM evaluators, even when provided with supervised data. We discover that misalignment in evaluation is not primarily due to biased priors over evaluation scores, but rather from misaligned evaluation standards, i.e., the likelihood of LLM evaluators defined in Equation (1). Therefore, we are motivated to explore a new evaluation paradigm with LLMs to facilitate a more aligned judgement.
@@
Figure 1: Illustration of RLHF (left), direct scoring with LLM evaluators (middle), and pairwise preference search (PairS; right). Pairwise preference data have been utilized to train the reward model to align the LLM in RLHF. Leveraging this idea, PairS reframes the traditional scoring-based evaluations into a series of pairwise comparisons between selected candidate pairs which are then used to form the final ranking.
@@
Inspired by how LLMs are aligned through a reward model with preference data in RLHF, we argue that LLM evaluators can produce more aligned and inherent evaluations by leveraging their pairwise ranking abilities. Recent work has started exploring this intrinsic property by prompting LLMs with a pair of candidates (Wang et al., 2023b; Chen et al., 2023; Qin et al., 2023; Li et al., 2024). However, the complexity and scalability of evaluating with pairwise preference have been largely overlooked. To produce a preference ranking for evaluation, existing methods require exhaustive pairwise comparisons, as they ignore the transitivity assumption [^1] The term transitivity in this work refers to mathematical transitivity, not linguistic transitivity. It is a property stating that if A is preferred to B and B is preferred to C, then A is preferred to C. (Qin et al., 2023). This makes the evaluation procedure intractable as the number of comparisons grows quadratically (Zheng et al., 2024; Liusie et al., 2024).
@@
In this work, we shed new light on the full potential of pairwise preference in LLM evaluators and its alignment with human judgements. We propose Pairwise-preference Search (PairS), an uncertainty-guided search method that efficiently estimates the Maximum Likelihood Estimate (MLE) of preference rankings by searching through an uncertainty-pruned pairwise comparison space. PairS is more effective and query-efficient than existing pairwise baselines while substantially outperforming traditional direct scoring evaluations. We extensively examine the applicability of PairS in representative evaluation tasks, including summarization and open-ended generation. We show that PairS produces more robust and transitive evaluations compared with the greedy search baseline. Lastly, we provide further insights into how pairwise preference can be used to quantify the transitivity ability of LLM evaluators and how it benefits from calibration.
@@
In sum, we provide the following contributions: 1) We present a systematic analysis of the limitations of calibration in aligning direct scoring LLM evaluators with human judgements. 2) We formulate the evaluation as ranking from the perspective of transitivity and propose PairS, an uncertainty-guided pairwise preference search method that efficiently estimates MLE rankings. 3) We demonstrate that PairS attains unique scalability in aligning LLM evaluations. We also provide insights into the transitivity and calibration of pairwise preferences in LLM evaluators.
@@ 2 On the Limitations of Calibration in Aligning LLM Evaluators
Figure 2: LLM evaluations are misaligned with human judgements. The score histograms on evaluating the coherence in HANNA (Chhun et al., 2022) and SummEval (Fabbri et al., 2021). We present the scores from gold human evaluations, LLMs, and LLMs after calibrations. The histograms can be interpreted as estimated score prior distributions via marginalization.
@@
where $x$ and $y$ form an input-output pair (e.g., a source text and its summary, respectively), $I$ denotes an instruction prompt, and $s$ represents the model’s rating score in its evaluation. Similarly, we can express the human posterior predictive distribution as $p_{\mathcal{H}}(s|y;x,I)$. For simplicity, we omit the $x$ and $I$ terms hereafter. When utilizing LLMs to perform Likert scale evaluations (Likert, 1932), it is known that there exists a mismatch between the score distributions of LLM ratings and human annotations (Liu et al., 2023b). We start by observing and interpreting the same misalignment phenomenon, as shown in Figure 2.
Since the posterior distribution consists of two components, the likelihood, and the score preference prior, we hypothesise that the discrepancy between the posterior distributions of LLMs and humans mainly arises from a misalignment between their respective likelihoods. To verify this hypothesis, we conduct a preliminary experiment by calibrating the LLM’s score preference prior, $p_{\mathcal{M}}(s)$ to match the human prior, $p_{\mathcal{H}}(s)$, described in what follows.
Score-Based LLM Evaluator Calibration. Following previous calibration methods (Zhou et al., 2024b; Jiang et al., 2023b), we first assume that the likelihood terms for both the LLM and humans are the same, $p_{\mathcal{M}}(y|s)=p_{\mathcal{H}}(y|s)$. Upon this assumption, the human posterior distribution $p_{\mathcal{H}}(s|y)$ can be expressed as:
@@
We can estimate LLM’s preference prior $p_{\mathcal{M}}(s)$ by marginalizing the model’s predictions, i.e., $p_{\mathcal{M}}(s)=\sum_{y}p_{\mathcal{M}}(s|y)p(y)$. Unlike previous work that assumes a uniform prior for $p_{\mathcal{H}}(s)$ (Zhou et al., 2024b), we directly estimate it by marginalizing human annotations to achieve a more accurate estimation. The calibration is performed by scaling the model’s predictions according to Equation (2).
@@ 3.1 Evaluation as Ranking
$\mathcal{L}_{NT}(\pi)=\prod_{i,j:\pi_{i}<\pi_{j}}P(y_{\pi_{i}}\succ y_{\pi_{j}}),$ (3)
where $\mathcal{L}_{NT}$ stands for the non-transitive likelihood, [^2] This likelihood can be derived from the general learning to rank likelihood in Appendix §C.1. which does not make any transitive assumptions for the pairwise comparison probabilities. The MLE ranking of $\mathcal{L}_{NT}$ is also equivalent to the Kemeny-Optimal Ranking (Kemeny, 1959). However, finding the optimal ranking is challenging in two aspects:
@@
However, strict transitivity is a strong assumption. To relax the strict transitivity assumption to some extent and find a better trade-off between computational complexity and performance, we propose simplifying the non-transitive likelihood in Equation (3) by assuming compositional stochastic transitivity (Latta, 1979; Oliveira et al., 2018). The derivation of this simplified likelihood is provided in Appendix §C.2:
@@
Figure 4: Illustration of PairS with a beam size of 2. Heatmap values show $P(A_{i}\succ B_{j})$ by Mistral 7B, deciding the search direction. Yellow-bordered lines highlight comparisons whose uncertainties are above $U_{h}$. Solid lines represent greedy search paths, while dashed lines denote alternative paths by other beam candidates.
3.2 Uncertainty-Guided Pairwise-Preference Search
@@
The PairS algorithm is described in Algorithm 1, and a demonstration of its workings is provided in Figure 3 and Figure 4. Intuitively, while the PairS-greedy chooses the preferred items between the heads of two sub-arrays in a greedy fashion, PairS-beam maintains a likelihood $\mathcal{L}_{T}$ beam that keeps track of the top-k trajectory candidates. By also considering less preferred items between the two heads, PairS explores a broader search space and potentially finds a better ranking.
@@
To represent the uncertainty between two items, we follow a commonly used approach in previous works (Wan et al., 2023a; b) that employs entropy as the uncertainty measurement. Hence, the uncertainty between a pair of comparisons can be expressed below:
@@ 3.3 A Scaling Variant of PairS
In the first stage, we sample and rank a subset of the evaluation set, referred to as the anchor set. The size of the anchor set can be statistically determined by accommodating a tolerable error margin (Cochran, 1977) (Further details and empirical evidence in Appendix §F.3). In the second stage, we use binary search to find the positions of the remaining samples relative to the anchor set, efficiently determining their appropriate positions in the overall ranking. As depicted in Figure 5, the second stage of the scaling method results in $O(N)$ complexity, a significant improvement over the original complexity. This two-stage scaling approach allows PairS to efficiently handle large evaluation sets while maintaining the benefits of beam search and uncertainty-based pruning. The scaling variant strikes a balance between complexity and ranking performance, making it suitable for real-world applications involving large-scale datasets.
@@ 3.4 On the Benefits of Pairwise-Preference Search
Transitivity. Both PairS-direct and PairS-beam assume transitivity to some extent, implying that an LLM with better transitivity can perform more accurate and more efficient evaluations. In other words, we expect better-performing LLMs to exhibit more transitive behaviour when making pairwise comparisons. Consequently, we propose that transitivity can serve as an effective measure of an LLM’s ability as an evaluator. By quantifying and comparing the transitivity of different LLMs, we can gain insight into their relative performance and suitability for evaluation tasks. To validate this hypothesis, we provide complementary experiments and analysis in Section 5.4.
@@
Calibration. Although pairwise preference inherently aligns LLM evaluations with human judgement, pairwise prompting can still suffer from misaligned preference priors, e.g., positional bias and contextual bias (Zhao et al., 2021; Zhou et al., 2024b). Therefore, we propose applying calibration on top of PairS to ensure debiased PairS evaluations. It is worth noting that, unlike calibrating the score-based LLM evaluators, which requires external information on the distribution of human preference prior, as shown in Section 2, pairwise comparisons can fully leverage a uniform human prior, due to the random nature of pairwise sample selections.
4 Related Work
@@
LLMs as Evaluators. Recently, LLMs have been increasingly used to automatically evaluate various natural language generation tasks (Zhong et al., 2022; Chen et al., 2023; Wang et al., 2023a; Tan et al., 2024). LLM evaluators enable reference-free or reference-guided metrics while being cheaper alternates to human evaluation in a diverse task set, such as summarization (Shen et al., 2023), machine translation (Kocmi & Federmann, 2023), open-ended story generation (Chiang & Lee, 2023), dialog response generation (Liu et al., 2024a) and instruction-following (Zeng et al., 2024). Unlike traditional reference-based evaluation metrics (Papineni et al., 2002; Liu et al., 2024b), LLM evaluators are usually equipped with evaluation criteria prompts and generate scores to access the quality of the candidate text (Liu et al., 2023a). This forms the scoring-based LLM evaluator that enables easy adaption to various tasks by simple instructions (Fu et al., 2023). To fairly examine various LLM evaluation abilities, efforts have also been devoted to developing meta-evaluation benchmarks (Wang et al., 2023b; Zhang et al., 2023; Zheng et al., 2024) and prompt optimization (Zhou et al., 2024a).
@@
Pairwise Preference. Preference data have been widely used in training the reward model to reflect human preferences in RLHF (Ouyang et al., 2022). For example, Llama-2 has leveraged binary preference with ‘chosen’ or ‘rejected’ responses in its reward modelling (Touvron et al., 2023). Following the preference idea, using pairwise comparisons in evaluating the LLM outputs has also been studied to some extent recently (Chen et al., 2023; Shen et al., 2023), where Wang et al. (2023b) find that pairwise comparison exhibits better human correlations than traditional scoring-based LLM evaluators. However, Zheng et al. (2024) point out the intractability of pairwise comparisons due to its quadratic query complexity (Liusie et al., 2024), which significantly limits its practicability. Qin et al. (2023) propose pairwise ranking with sorting but ignore the important transitivity of LLM predictions. In our work, we address the aforementioned issues by searching in a tractable pairwise search and transitive space based on model uncertainty.